OpenStudy (anonymous):

what is the name of the shape graphed by the function r=2cos(3 theta)? A. rose with 3 pedals B. limacon with inter loop C. line D. circle

OpenStudy (anonymous):

A

OpenStudy (rogue):

Limacons follow the format:\[r = a + b \cos \theta\]

OpenStudy (rogue):

Rose follows:\[r = a \sin n \theta\]\[r = a \cos n \theta\]

OpenStudy (rogue):

So its choice A.